OpenStudy (anonymous):

show in steps thank you 3/4(x+2)=-1

OpenStudy (anonymous):

(3/4)(x+2)=-1 ===> multiply both sid in (4/3) ===> x+2=-4/3 ===> subtract 2 from both side ==> x=(-4/3) -2 =(-4/3) -(6/3) =-10/3

OpenStudy (anonymous):

\[x+2=-\frac{4}{3}\] \[x=-\frac{4}{3}-2\] \[x=-\frac{10}{3}\]

OpenStudy (anonymous):

thank u so much

OpenStudy (anonymous):

yw